EXPERIMENTAL AND THEORETICAL STUDIES OF Mn_{2}(CO)_{10}

\underline{HOLLY B. LAVENDER}, TERRY GUSTAFSON, BRUCE BURSTEN, AND MARCELLO VITALE, Department of Chemistry, The Ohio State University, Columbus, Ohio 43210.

Ground state resonance Raman spectroscopy, Fenske-Hall molecular orbital calculations, and density functional theory were utilized to study the photochemistry of Mn_{2}(CO)_{10} in an attemt to understand better the photochemical reaction pathways of dinuclear transition metal organometallic complexes. In this work, resonance enhancement of the e_{3} Mn-C-O bending mode, the e_{3} C-O stretching vibration and the a_{1} Mn-Mn bond stretch are found to play an important role in the resonance Raman spectra of Mn_{2}(CO)_{10}. These data, together with the theoretical calculations, are consistent with the proposal that the photochemical intermediate Mn_{2}(CO)_{9} contains a linear semi-bridging CO.